# FAST functional connectivity implicates P300 connectivity in working memory deficits in Alzheimer’s disease

### v1

@@ Abstract
Measuring transient functional connectivity is an important challenge in Electroencephalogram (EEG) research. Here, the rich potential for insightful, discriminative information of brain activity offered by high temporal resolution is confounded by the inherent noise of the medium and the spurious nature of correlations computed over short temporal windows. We propose a novel methodology to overcome these problems called Filter Average Short-Term (FAST) functional connectivity. First, long-term, stable, functional connectivity is averaged across an entire study cohort for a given pair of Visual Short Term Memory (VSTM) tasks. The resulting average connectivity matrix, containing information on the strongest general connections for the tasks, is used as a filter to analyse the transient high-temporal resolution functional connectivity of individual subjects. In simulations, we show that this method accurately discriminates differences in noisy Event-Related Potentials (ERPs) between two conditions where standard connectivity and other comparable methods fail. We then apply this to analyse activity related to visual short-term memory binding deficits in two cohorts of familial and sporadic Alzheimer’s disease. Reproducible significant differences were found in the binding task with no significant difference in the shape task in the P300 ERP range. This allows new sensitive measurements of transient functional connectivity, which can be implemented to obtain results of clinical significance.
I Introduction
@@
Network Science approaches to the analysis of complex networks provide useful tools for the analysis of connectivity between agents [1, 2] . The brain is an example of a complex network where pair-wise dependencies between brain regions are of value in the detection of cognitive phenomena. It is found that it is neither spatial nor temporal localization of brain activity that underpins cognitive phenomena and the corresponding brain function but in fact, how the different areas of the brain are dynamically interconnected over time [3, 4]. This has led to a boom in studies of functional connectivity of brain activity across viable formats– mainly the blood oxygenation level-dependent signal in MRI and electromagnetic recordings from EEG and MEG. Here, typically, signals from parcellated regions (in fMRI) or sensors (in EEG/MEG) are subject to pairwise measures of connectivity, such as correlation coefficients, coherence measures, or phase-based measures.
@@
The electroencephalogram (EEG) contains important discriminating information relating to sequential brain processes in response to various cognitive tasks. Providing a very high temporal resolution, scalp EEG allows for the direct recording of electromagnetic activity of the brain in a non-invasive, relatively cheap way [5]. Scalp EEG presents several notable limitations however, with the most prominent being the substantial noise levels inherent in the recorded signals. This noise poses a significant challenge, especially when attempting to investigate the functional connectivity associated with transient cognitive processes occurring within brief time-frames, typically spanning mere tens of milliseconds. A pivotal issue within the realm of functional connectivity of EEG signals pertains to the extraction of dependable connectivity estimates within these remarkably short time intervals [6]. This problem underscores the necessity for novel methodologies to overcome noise-related hurdles and facilitate the precise examination of cognitive processes unfolding at rapid temporal scales [1]. Measuring dynamic functional brain connectivity in short time windows is gaining increasing recognition in AD research due to its potential to provide information for the early detection of the devastating disease [7, 8]. An important reason for this is the growing recognition that intricate changes in brain connectivity can occur before the onset of clinical symptoms; this makes it a promising avenue for early biomarker development and a better understanding of disease progression. AD is not a static condition but involves dynamic changes in brain function, short-time based analysis with non-invasive brain imaging techniques can provide important breakthroughs in AD early detection, especially in low-income countries [9].
@@
Despite the growing popularity of these studies, there has been limited methodological work on the analysis of EEG dynamic functional connectivity (DFC). Previous work typically focuses on the sliding window method [10, 11, 12], while this is fairly effective, the temporal resolution and susceptibility to noise is largely determined by the window size [27]. It has become a priority to simultaneously improve the temporal resolution of DFC while being robust to spurious connections and noise. Methods such as the Short-Term Fourier Transform [11] and Wavelet Analysis [13] have been frequently applied in this domain, but once again, the dependency on window size causes bottlenecks in regard to temporal resolution and noise robustness.
@@
Graph Signal Processing (GSP) approaches have been employed frequently in the past to perform spectral analysis of signals in the graph domain, as opposed to the temporal domain [14]. This is achieved by computing the eigenvalue decomposition of a relevant Graph-Shift operator such as the Graph Laplacian or Adjacency matrix followed by the Graph Fourier Transform. However, the frequencies that emerge through the graph eigenvectors are still determined completely by the graph topology and do not involve the signal [1].
@@
Here, we propose a new method for extracting reliable estimates of short-term functional connectivity. This is based on Graph Variate Signal Analysis [1], a more general framework for graph signals. Specifically, it describes how to leverage graphs of long-term reliable connectivity information to filter instantaneous bivariate node functions of multivariate signals. In essence, this emphasizes important connections and minimizes spurious ones (a well-known issue in EEG signals). This gives us a readily interpretable method to analyse the transient changes in brain activity at a high temporal resolution using pairwise connectivity measures (i.e. correlation) between EEG electrodes. Graph-variate dynamic connectivity [1] is when the long-term connectivity estimate is computed from the signal itself over the given epoch of interest so that the graph signal is directly related to the underlying graph and measurements and, therefore, solely relates to one connectivity function. Here, we develop and employ a novel methodology based on GVD connectivity which we call Filter Average Short-Term (FAST) connectivity. Essentially, FAST connectivity uses the average long-term connectivity matrix over the whole study cohort as a filter of transient functional connectivity at the individual level. Essentially, we are deriving the most consistent connections across all participants and then asking if the temporal activity associated with those connections shows differences between, for example, patients and control. As we shall see, the high temporal resolution of brain activity provided by the Electroencephalogram(EEG) [5] can now be exploited to detect more sensitive and specific cognitive changes in very short time frames.
@@
We demonstrate the power of our novel FAST Connectivity methodology in simulations for picking out the true activity of ERPs in the presence of different levels of noise and different numbers of trials. We then apply this to the dataset containing EEG signals from the participants in the Visual Short Term Memory (VSTM) tasks [15]. Following this we perform rigorous statistical testing on temporal windows resulting from the multi-layer graph-variate tensor. This uncovers a potential biomarker for the early detection of AD as well as a potential indicator of disease progression from the EEG signals. This is a first in the formal application of GVD connectivity [1] to EEG signals.
II Methods
II-A Background
The novel method proposed is a form of graph-variate signal analysis [1]. Graph-variate signal analysis is defined formally as:
@@ II-A Background
Where the formula defines the bi-variate analysis of the multivariate signal X filtered by the corresponding static matrix W of the graph-variate signal. $\underline{\mathbf{J}}_{(t)}$ denotes the t’th n x n matrix of $\underline{\mathbf{J}}$ and $\circ$ is the Hadamard product. Each timestep of $\underline{\mathbf{J}}$ is defined by a $n\times n$ matrix computed using the pairwise bi-variate connectivity values between signal pairs. The form of dynamic Connectivity is determined by the node function $F_{V}$. In our analysis, we used the squared difference inspired by the reformulation of the graph Dirichlet Energy [2] and also the instantaneous Pearson Correlation values as readily interpretable forms of connectivity.
@@
[…]
@@
$c_{ij}=\frac{\sum_{t\epsilon T}(x_{i}(t)-\bar{x_{i}})(x_{j}-\bar{x_{j}}))}{\sqrt{\sum_{t\epsilon T}(x_{i}(t)-\bar{x_{i}})^{2}}\sqrt{\sum_{t}(x_{j}(t)-\bar{x_{j}})^{2}}}$ (3)
@@
combining this with the squared difference or instantaneous correlation as the node function, GVD connectivity is defined as:
@@
using the squared difference node function, and
$\theta({\textbf{x}}_{i},{\textbf{x}}_{j},t)=\rho_{t}(i,j)=c_{ij}\left|(x_{i}(t)-\bar{\textbf{x}_{i}})(x_{j}(t)-\bar{\textbf{x}_{j}})\right|,$ (5)
using the Instantaneous correlation node function.
This can be considered as a representation of the amplitude and, thus, an analysis in this domain.
While we can average over a subset of nodes for a modular comparison between regions of the brain, we are more focused on a Global analysis in this paper.
II-B FAST Connectivity
We now present FAST Connectivity, a novel form of Graph Variate Signal Analysis. We propose a singular filter for all participants in time-locked cognitive task-based experiments. The filter takes the long-term connectivity estimates of all participants in the experiments and averages over them to create a single FAST filter for all participants that automatically emphasizes important connections and suppresses spurious ones in the general time-locked cognitive task of interest (in this case the VSTM binding and shape tasks). We define the FAST filter as:
@@ II-B FAST Connectivity
Where $C$ is the matrix of the absolute values of the individual long-term correlation estimates, with $c_{ij}$ representing Each entry in the matrix. For $P=1,2…N$, where $P$ is each participant and N is the total number of participants. We define our FAST filter as:
@@
$c_{ij}=\left|\frac{\sum_{t\epsilon T}(x_{i}(t)-\bar{x_{i}})(x_{j}-\bar{x_{j}}))}{\sqrt{\sum_{t\epsilon T}(x_{i}(t)-\bar{x_{i}})^{2}}\sqrt{\sum_{t}(x_{j}(t)-\bar{x_{j}})^{2}}}\right|$ (7)
[1]
@@
For each $P=1…N$ where $N$ is the total number of participants the same FAST filter is applied to each participant. The squared difference is our node function of choice.
@@
FAST connectivity proposes the same filter for all participants $P$. This filter is constructed using the magnitude of the stable long-term correlation averaged over all participants. Essentially FAST connectivity can be defined as a Graph-Variate Signal Analysis where the long-term stable matrix is the FAST Filter and the node function is the squared difference inspired by the Dirichlet Energy.
The squared difference between signal pair values can be considered as a localized measure of the variation between signal pair values. This captures the local variation of the signal. A higher value would indicate higher variation in the signal pair region, whereas if it was small, the signal pair values are fairly constant or change smoothly in the localized region. It is important to stress that due to the nature of the squared difference node function, it emphasises anti-correlative information as higher squared difference values (that usually indicate a strong anti-correlative interaction between signal pairs) in the instantaneous connectivity profiles contribute most to the value of the mean network metrics.
On the other hand, Filtering the instantaneous correlation values with the long-term correlation coefficient of all participants gives more weight to regions where the local variation aligns with the long-term correlation patterns across participants allowing for a selective emphasis on specific features that are both locally and globally consistent. Overall, The FAST connectivity analysis is sensitive to both fine-scale variations within individual EEG signals and broader patterns shared across participants. By combining a global measure with local information, the method is effective in identifying regions that not only vary locally but also exhibit synchronized variations across participants. This should reflect meaningful functional connectivity patterns while reducing noise and spurious connectivity.
@@
This approach can enhance the detection of relevant ERPs, especially in cases where the direction of changes may vary or fluctuate between negative and positive values, providing a robust and direction-agnostic measure for identifying meaningful features in small temporal windows.
Overall, it is useful to think of the FAST filter as a ‘rating’ mechanism, essentially telling us which nodes are of particular interest when analyzing individual dynamic functional connectivity. This, combined with the squared difference node function, allows us to determine areas of high variation between ‘important’ graph signal pairs.
As mentioned before, similar to the Modular Dirichlet Energy (MDE) [2] a prototype of GVD Connectivity, FAST connectivity analyzes the temporal brain networks from a unique angle compared to other approaches such as time series analysis of network metrics. Essentially, one stable network of long-term connectivity is computed over the whole epoch, which is used as a support for localized analysis of very small temporal windows, allowing us to maximize the high temporal resolution of EEG signals. The activity is, in fact, encoded in the graph signal itself rather than the time-varying edge weights. This allows for analysis of smaller temporal windows of activity and also analysis of the overall long-term activity.
@@
II-C Network analysis of FAST connectivity

### v2

@@ Abstract
Measuring transient functional connectivity is an important challenge in Electroencephalogram (EEG) research. Here, the rich potential for insightful, discriminative information of brain activity offered by high temporal resolution is confounded by the inherent noise of the medium and the spurious nature of correlations computed over short temporal windows. We propose a methodology to overcome these problems called Filter Average Short-Term (FAST) functional connectivity. First, long-term, stable, functional connectivity is averaged across an entire study cohort for a given pair of Visual Short Term Memory (VSTM) tasks. The resulting average connectivity matrix, containing information on the strongest general connections for the tasks, is used as a filter to analyse the transient high-temporal resolution functional connectivity of individual subjects. In simulations, we show that this method accurately discriminates differences in noisy Event-Related Potentials (ERPs) between two conditions where standard connectivity and other comparable methods fail. We then apply this to analyse activity related to visual short-term memory binding deficits in two cohorts of familial and sporadic Alzheimer’s Disease(AD) related Mild Cognitive Impairment(MCI). Reproducible significant differences were found in the binding task with no significant difference in the shape task in the P300 ERP range. This allows new sensitive measurements of transient functional connectivity, which can be implemented to obtain results of clinical significance.
keywords:
Alzheimer’s Disease, , P300, Working memory, Dynamic Functional Connectivity, EEG, FAST
articletype: Research
Computer and Information Sciences, University of Strathclyde, Glasgow, UK Latin American Brain Health Institute (BrainLat), Universidad Adolfo Ibañez, Santiago, Chile Neuroscience Group of Antioquia, Medicine School, University of Antioquia, Medellín, Colombia Global Brain Health Institute, Trinity College Dublin, Ireland Psychological Sciences and Health, University of Strathclyde, Glasgow, UK
[…]
Om Royo.roy.2022@uni.strath.ac.uk
{authorsummary}
FAST connectivity is an EEG analysis method that enhances detection of dynamic functional connectivity changes during cognitive events like ERPs, effectively handling EEG noise and maximizing temporal resolution. It reduces required trial numbers for reliable analysis, particularly beneficial for studying tasks such as working memory. FAST connectivity complements traditional methods by focusing on temporal connectivity patterns, showing superior performance in simulations compared to standard measures. Applied to Alzheimer’s datasets, it identifies significant differences in brain activity during visual short-term memory tasks, highlighting its potential for understanding neurological conditions.
1 Introduction
@@
Network Science approaches to the analysis of complex networks provide useful tools for the analysis of connectivity between agents [1, 2] . The brain is an example of a complex network where pair-wise dependencies between brain regions are of value in the detection of cognitive phenomena. It is found that it is neither spatial nor temporal localization of brain activity that underpins cognitive phenomena and the corresponding brain function but in fact, how the different areas of the brain are dynamically interconnected over time [3, 4, 5]. This has led to a boom in studies of functional connectivity of brain activity across viable formats[7, 6, 8] – mainly the blood oxygenation level-dependent signal in fMRI[9] and electromagnetic recordings from EEG and MEG. Here, typically, signals from parcellated regions (in fMRI) or sensors (in EEG/MEG) are subject to pairwise measures of connectivity, such as correlation coefficients, coherence measures, or phase-based measures[10, 11] . In particular there has been a clear increase in the study of functional connectivity changes related to Alzheimer’s Disease (AD) in the form of the study of the brain as a network using various graph-theoretic tools [12, 13, 14]
@@
The electroencephalogram (EEG) contains important discriminating information relating to sequential brain processes in response to various cognitive tasks[17, 16, 15]. Providing a very high temporal resolution, scalp EEG allows for the direct recording of electromagnetic activity of the brain in a non-invasive, relatively cheap way [18]. Scalp EEG presents several notable limitations however, with the most prominent being the substantial noise levels inherent in the recorded signals. This noise poses a significant challenge, especially when attempting to investigate the functional connectivity associated with transient cognitive processes occurring within brief time-frames, typically spanning mere tens of milliseconds. A pivotal issue within the realm of functional connectivity of EEG signals pertains to the extraction of dependable connectivity estimates within these remarkably short time intervals [19]. This problem underscores the necessity for novel methodologies to overcome noise-related hurdles and facilitate the precise examination of cognitive processes unfolding at rapid temporal scales [1]. Measuring dynamic functional brain connectivity in short time windows is gaining increasing recognition in AD research due to its potential to provide information for the early detection of the devastating disease [20, 21, 34]. An important reason for this is the growing recognition that intricate changes in brain connectivity can occur before the onset of clinical symptoms; this makes it a promising avenue for early biomarker development and a better understanding of disease progression. AD is not a static condition but involves dynamic changes in brain function, short-time based analysis with non-invasive brain imaging techniques can provide important breakthroughs in AD early detection, especially in low-income countries [22].
@@
Despite the growing popularity of these studies, there has been limited methodological work on the analysis of EEG dynamic functional connectivity (DFC). Previous work typically focuses on the sliding window method [23, 24, 25], while this is fairly effective, the temporal resolution and susceptibility to noise is largely determined by the window size. It has become a priority to simultaneously improve the temporal resolution of DFC while being robust to spurious connections and noise. Methods such as the Short-Term Fourier Transform [24] and Wavelet Analysis [26] have been frequently applied in this domain, but once again, the dependency on window size causes bottlenecks in regard to temporal resolution and noise robustness.
@@
Graph Signal Processing (GSP) approaches have been employed frequently in the past to perform spectral analysis of signals in the graph domain[27]. This is achieved by computing the eigenvalue decomposition of a relevant Graph-Shift operator such as the Graph Laplacian or Adjacency matrix followed by the Graph Fourier Transform. However, the frequencies that emerge through the graph eigenvectors are still determined completely by the graph topology and do not involve the signal itself [1].
@@
Here, we propose a new method for extracting reliable estimates of short-term functional connectivity. This is based on Graph Variate Signal Analysis [1], a more general framework for graph signals. Specifically, it describes how to leverage graphs of long-term reliable connectivity information to filter instantaneous bivariate node functions of multivariate signals. In essence, this emphasizes important connections and minimizes spurious ones (a well-known issue in EEG signals). This gives us a readily interpretable method to analyse the transient changes in brain activity at a high temporal resolution using pairwise connectivity measures between EEG electrodes.
Graph-variate dynamic connectivity [1] is when the long-term connectivity estimate is computed from the signal itself over the given epoch of interest so that the graph signal is directly related to the underlying graph and measurements and, therefore, solely relates to one connectivity function.
Here, we develop and employ a novel methodology based on GVD connectivity which we call Filter Average Short-Term (FAST) connectivity. Essentially, FAST connectivity uses the average long-term connectivity matrix over the whole study cohort as a filter of transient functional connectivity at the individual level. Essentially, we are deriving the most consistent connections across all participants and then asking if the temporal activity associated with those connections shows differences between, for example, patients and control.
Traditional functional connectivity methods typically employ measures such as the amplitude envelope correlation (AEC) or phase locking value (PLV) combined with source reconstruction methods to assess pairwise functional coupling[29]. However, the inherent noise in EEG recordings undermines the efficacy of using Pearson correlation coefficients between channel time series as a reliable measure. The FAST filter, which provides a noise-robust matrix representing consistent long-term correlations as a stable support for instantaneous connectivity rather than being the conclusive object of analysis, offers a more appropriate use of this metric for analyzing temporally evolving instantaneous connectivity. This enhances the reliability of coupling measures in the presence of EEG noise.
EEG recordings are further complicated by individual variability, heterogeneous artifacts, volume conduction effects, and low spatial resolution, which pose significant challenges for spatial filtering approaches such as source reconstruction methods. These methods struggle to accurately solve the inverse problem of mapping scalp-recorded activity to specific brain regions, especially given the unknown number of sources at any given time. FAST connectivity does not operate in this more traditional domain of analysis and is not intended to replace spatial methods for identifying precise brain regions of activity. Instead, it generalizes consistent statistical dependencies across broader brain regions and serves as a temporal filtering technique. This makes it a valuable complement to spatial filtering approaches, enhancing the overall analysis of brain connectivity.
As we shall see, the high temporal resolution of brain activity provided by the Electroencephalogram(EEG) [18] can now be exploited to detect more sensitive and specific cognitive changes in very short time frames.
@@
We demonstrate the power of FAST Connectivity in simulations for picking out the true activity of ERPs in the presence of different levels of noise and different numbers of trials. We then apply this to the dataset containing EEG signals from the participants in the Visual Short Term Memory (VSTM) tasks [28]. Following this we perform rigorous statistical testing on temporal windows resulting from the multi-layer graph-variate tensor. This uncovers a potential biomarker for the early detection of AD.
2 Methods
2.1 Background
The method proposed is inspired from the modular dirichlet energy [2] and graph-variate signal analysis [1] methods. We thus briefly introduce these concepts.
The Dirichlet energy of a graph signal $\mathbf{x}$ is defined as:
$\mathbf{x}^{T}L\mathbf{x}=\sum_{i,j=1}^{n}w_{ij}(x_{i}-x_{j})^{2}$ (1)
[2].
Essentially this allows us to contrast pair-wise graph signal smoothness or variability with a measure $w_{ij}$.
The squared difference between signal pair values can be considered as a localized measure of the variation between signal pair values. This captures the local variation of the signal. A higher value would indicate higher variation in the signal pair region, whereas if it was small, the signal pair values are fairly constant or change smoothly in the localized region. The Dirichlet energy captures the sum of the local variations over the graph. The term Local Dirichlet energy will be used from now to refer to pairwise or modular squared differences.
Graph-variate signal analysis is defined formally as:
@@ 2.1 Background
Where the formula defines the bi-variate analysis of the multivariate signal X filtered by the corresponding static matrix W of the graph-variate signal. $\underline{\mathbf{J}}_{(t)}$ denotes the t’th n x n matrix of $\underline{\mathbf{J}}$ and $\circ$ is the Hadamard product. Each timestep of $\underline{\mathbf{J}}$ is defined by a $n\times n$ matrix computed using the pairwise bi-variate connectivity values between signal pairs. The form of dynamic Connectivity is determined by the node function $F_{V}$.
@@
$\theta(\mathbf{x}_{\textit{i}},\mathbf{x}_{\textit{j}},\textit{t})=\begin{cases}c_{ij}F_{v}(x_{i}(t),x_{j}(t)),&\text{if }i\neq j\\ 0,&\text{if }i=j\end{cases}$ (3)
@@
$c_{ij}=\frac{\sum_{t\in T}(x_{i}(t)-\mathbf{\bar{x}}_{i})(x_{j}(t)-\mathbf{\bar{x}}_{j})}{\sqrt{\sum_{t\in T}(x_{i}(t)-\mathbf{\bar{x}}_{i})^{2}}\sqrt{\sum_{t\in T}(x_{j}(t)-\mathbf{\bar{x}}_{j})^{2}}},$ (4)
@@
where $T$ is the epoch of interest and $\mathbf{\bar{x}}_{i}$ is the mean of the values over time of the node $i$.
combining this with the squared difference, GVD connectivity can be defined as:
@@
Where $\tilde{x}_{i}(t)$ is the normalised signal over the node space:
$\tilde{x}_{i}(t)=\frac{x_{i}(t)-\mathbf{\bar{x}}(t)}{\sqrt{\frac{1}{n-1}\sum_{k=1}^{n}\left(x_{k}(t)-\mathbf{\bar{x}}(t)\right)^{2}}},$ (6)
and $\mathbf{\bar{x}}(t)$ is the mean over nodes of the signal at time $t$:
$\mathbf{\bar{x}}(t)=\frac{1}{n}\sum_{k=1}^{n}x_{k}(t).$ (7)
It is clear now that the reformulated Dirichlet energy is a special case of Graph-Variate Signal Analysis.
2.2 FAST Connectivity
We now present FAST Connectivity. A single filter is proposed for all participants in time-locked cognitive task-based experiments. The filter takes the long-term connectivity estimates of all participants in the experiments and averages over them to create a single FAST filter for all participants that automatically emphasizes important connections and suppresses spurious ones in the general time-locked cognitive task of interest (in this case the VSTM binding and shape tasks). We define the FAST filter as:
@@ 2.2 FAST Connectivity
[…]
@@
$c_{ij}=\left|\frac{\sum_{t\in T}(x_{i}(t)-\mathbf{\bar{x}}_{i})(x_{j}(t)-\mathbf{\bar{x}}_{j})}{\sqrt{\sum_{t\in T}(x_{i}(t)-\mathbf{\bar{x}}_{i})^{2}}\sqrt{\sum_{t\in T}(x_{j}(t)-\mathbf{\bar{x}}_{j})^{2}}}\right|,$ (9)
@@
$\theta^{\text{FAST}}(\mathbf{x}_{\textit{i}}^{P},\mathbf{x}_{\textit{j}}^{P},\textit{t})=\begin{cases}c_{ij}^{\text{FAST}}(\tilde{x}_{i}^{P}(t)-\tilde{x}_{j}^{P}(t))^{2},&\text{if }i\neq j\\ 0,&\text{if }i=j\end{cases}$ (10)
FAST connectivity proposes the same filter for all participants $P$. This filter is constructed using the magnitude of the stable long-term correlation averaged over all participants.
Setting $w_{ij}$ as the relevant entry of the FAST filter matrix allows us to weight corresponding instantaneous variability by long-term stable connections. In Layman terms, we first determine which brain regions are most consistently strongly connected in terms of statistical dependencies over the whole cohort , we are then ’focusing’ on these regions and analyzing the variability in these localized regions using the squared difference function.
Overall, The FAST connectivity analysis is sensitive to both fine-scale variations within individual EEG signals and broader patterns shared across participants. By combining a global measure with local information, the method is effective in identifying regions that not only vary locally but also exhibit strong synchronized connectivity across participants. This should reflect meaningful functional connectivity patterns while reducing noise and spurious connectivity.
@@
Similar to the Modular Dirichlet Energy (MDE) [2] a prototype of GVD connectivity, FAST connectivity analyzes the temporal brain networks from a unique angle compared to other approaches such as time series analysis of network metrics. Essentially, one stable network of long-term connectivity is computed over the whole epoch and averaged over all participants, this is used as a support for localized analysis of very small temporal windows, allowing us to maximize the high temporal resolution of EEG signals. The activity is, in fact, encoded in the graph signal itself on the temporally evolving edge weights. This allows for analysis of smaller temporal windows of activity and also analysis of the overall long-term activity.
@@
2.3 Network analysis of FAST connectivity